OpenStudy (itrymath):

Meg initially has 3 hours of pop music and 2 hours of classical music in her collection. Every month onwards, the hours of pop music in her collection is 5% more than what she had the previous month. Her classical music does not change. Which function shows the total hours of music she will have in her collection after x months? (4 points) f(x) = 2(0.05) x + 3 f(x) = 2(0.05) x + 3 f(x) = 2(1.05) x+ 3 f(x) = 3(1.05) x + 2

OpenStudy (itrymath):

@YanaSidlinskiy @mathmale @ShadowLegendX

OpenStudy (itrymath):

D can be eliminated

OpenStudy (yanasidlinskiy):

No.... It's the right answer.

OpenStudy (itrymath):

:)

OpenStudy (itrymath):

jk

OpenStudy (itrymath):

answer is C

jimthompson5910 (jim_thompson5910):

@YanaSidlinskiy is correct x = 0 is the starting month plug this into the function to get f(x) = 3(1.05)^x + 2 f(0) = 3(1.05)^0 + 2 f(0) = 3(1) + 2 f(0) = 3 + 2 f(0) = 5 so there's 5 hours total as expected in the beginning ----------------------------------------------- One month later, there's 5 more percent of pop music. So take 5% of 3 hours to get 0.05*3 = 0.15 which gets added on to the previous total 5+0.15 = 5.15 if we plug x = 1 into the function, we should get 5.15. Let's find out f(x) = 3(1.05)^x + 2 f(1) = 3(1.05)^1 + 2 f(1) = 3(1.05) + 2 f(1) = 3.15 + 2 f(1) = 5.15 So it fits. This pattern would keep going. I recommend you trying out the next step to see if it works or not

OpenStudy (itrymath):

@jim_thompson5910 its D isnt it

jimthompson5910 (jim_thompson5910):

Where are you stuck?

OpenStudy (itrymath):

im not

jimthompson5910 (jim_thompson5910):

yes the answer is D

OpenStudy (itrymath):

Every month onwards, the hours of pop music in her collection is 5% more than what she had the previous month.

OpenStudy (itrymath):

this says it all

OpenStudy (itrymath):

@YanaSidlinskiy

OpenStudy (itrymath):

you were correct

OpenStudy (yanasidlinskiy):

Thank you:) Now if you could stop with the jokes and knock it off I'd appreciate it.